OpenStudy (anonymous):

3x-2y=5 , 4x-5y= -26 ( clear the system of fractions and then solve by substitution)

OpenStudy (anonymous):

Okay first you set them up like this 3x-2y=5 4x-5y=-26 Then you make one of the variable 0 Okay let's make y zero so we can solve for x. To make y 0 we have to multiply bot top and bottom so that they add to each other and become 0. (3x-2y=5)5 (4x-5y=-26)2 Can you try to solve it from here?